# Exploring baby Julia sets in parameter space slices for Generalized McMullen Maps

### v1

@@ 7. Topological Disks in the $r_{n,a}$ Parameter Space
Here we consider $k=2j,$ for $j\in\{1,\ldots,n-1\}$ (that is, $k$ even and $k\in\{2,\ldots,2n-2\}$).
First: for $a=a_{2j}$ (as defined in Lemma 5.1) we have $r(v_{+})=v_{+}$ and $r(v_{-})=v_{+}$. Further, $v_{-}$ itself is a critical point, with
@@ 7.1. Dynamical plane
Now, recall (Lemma 3.1) that $D_{+}$ (or $D_{+}(a)$ if this is needed for clarity) is the Fatou component containing $v_{+}$ (aka $v_{+}(a)$) and on $\overline{D_{+}},$ the map $r$ is conjugate to $p_{0}(z)=z^{2}$ via $\phi_{a}$; not just for $a=a_{2j}$, but for all $a\in\mathcal{H}=\{a|\ r_{n,a}$ is hyperbolic }\}.
@@
Assume for $a=a_{2j},$ $v_{-}\notin D_{+}$ that is $D_{-}\neq D_{+}$.
First we discuss the dynamics at $a=a_{2j}$. Now $v_{-}$ is not a fixed point for $a=a_{2j},$ and so there are $n$ critical points which map to $v_{-}$ and are all distinct from it. From Lemma 2.13, the critical points $w_{2\ell-1}=|a|^{1/2n}e^{i(\frac{\psi+2(2\ell-1)\pi}{2n})}$ for $\ell=1,2,3,...,n,$ map to $v_{-}$, and the critical points $w_{2\ell}=|a|^{1/2n}e^{i(\frac{\psi+2(2\ell)\pi}{2n})}$ for $\ell=0,1,2,...,n-1,$ map to $v_{+},$ with $w_{0}=v_{+}$ in this case and our special $j$ satisfying $w_{2j}=v_{-}$. So, not counting $v_{+}$ itself, there are n1n-1 critical points mapping to $v_{+},$ and $n$ critical points mapping to $v_{-}$ (and those $n$ mapping to $v_{-}$ don’t include $v_{-}$ since it maps to $v_{+})$.
For each $\ell\in\{0,\ldots,2n-1\},$ let $D_{\ell}$ be the Fatou component containing the critical point $w_{l}$, so $D_{0}=D_{+}$ and for $a=a_{2j}$, $D_{2j}=D_{-}.$
Thus $R$ maps each $D_{2\ell-1}$ onto $D_{-}$ as a $2$ to $1$ branched covering ramified over the critical point. Similarly, $R$ maps each $D_{2\ell},$ including $D_{-}$, as a $2:1$ branched cover onto $D_{+}$, ramified over the critical point in $D_{2\ell}$.
See Figure 10 for an illustration of how the Fatou components containing critical points map, in these cases of $a=a_{2j}$ so $v_{+}$ fixed and $v_{-}$ mapping onto $v_{+}$.
@@
Next, if we consider $a$ near $a_{2j}$, of course $D_{+}$ varies with $a$, but (see Lemma 3.1), $v_{+}$ is safely inside of $D_{+}$, hence for $a$ near $a_{2j},v_{-}$ is safely inside of $D_{-}$ (since $r(v_{-})=v_{+}$ at $a=a_{2j}$). As $a$ varies away from $a_{2j},v_{-}$ moves toward the boundary of $D_{-}$, its image $r(v_{-})$ moves away from $v_{-}$ (heading toward the boundary of $D_{+}$ as $a$ grows) and the critical point pre-images of $v_{-}$, the $w_{2\ell-1}$’s, approach the boundary of their Fatou components $D_{2\ell-1}$. Or, it’s more accurate to say the boundary of the $D_{2\ell-1}$’s tend to the $w_{2\ell-1}$’s. Since $R$ is a $2:1$ ramified cover from each component $D_{2\ell-1}$ of $R^{-1}(D_{-})$ onto $D_{-}$, as $v_{-}\to\partial D_{-}$, when $v_{-}$ hits the boundary of $D_{-}$ you then must have each $D_{2\ell-1}$ devolve into a topological lemniscate, with a figure 8 boundary with the critical point $w_{2\ell-1}$ at the crossing.
As $a$ continues to move away from $a_{2j},$ just past the lemniscate phase there are then two Fatou components that replace $D_{2\ell-1},$ each mapping $1:1$ (because they have lost their critical point) onto the preimage of $D_{+}$ that is no longer $D_{-}$ because $v_{-}$ has moved out of it.
@@
Now we apply this result in our case here where $\overline{D_{+}}$ is the baby filled Julia set $K_{+}$, and the quadratic polynomial we are conjugate to is $P_{0}(z)=z^{2}$ with filled Julia set $\overline{\mathbb{D}}$. Then we get a set of angle assignments on each preimage of $\partial D_{+}$, which respects the dynamics of $r_{n,a}$.
@@
Let $r=r_{n,a}$ be in the subfamily given in (1), so the Julia set $J(r)=J_{r}$ contains a baby Julia set $\overline{D_{+}}$ on which $r$ is conjugate to $P_{0}(z)=z^{2}$ on its Julia set, the closed unit disk (via a map $\phi_{a}$ which is q.c. in a neighborhood of the filled Julia sets and analytic on the interior). Assume in addition that $v_{-}$ lies in a preimage of $D_{+}$.
@@
    For components $D_{2\ell}$, $\ell=0,\dots,n-1$, each containing the critical point $w_{2\ell}$, (observe this includes $D_{+}=D_{0}$, and $D_{-}$, for some $\ell>0$), we have
    $r\Big(\Gamma|^{\partial D_{2\ell}}\big([0,\frac{1}{2})\big)\Big)=r\Big(\Gamma|^{\partial D_{2\ell}}\big([\frac{1}{2},1)\big)\Big)=\Gamma|^{\partial D_{+}}\big([0,1)\big)$.
@@
For (1), note $r(D_{2\ell})=D_{+}$ for each $\ell=0,\dots,n-1$, and $r(D_{2\ell})=D_{+}$ is a $2:1$ mapping.
@@
Then each of the statements about how subintervals $0,\frac{1}{2}]$, $\frac{1}{2},1]$, $0,\frac{1}{4}]$, $\frac{1}{4},\frac{1}{2}]$, $\frac{1}{2},\frac{3}{4}]$, and $\frac{3}{4},1]$ map follows directly from applying Theorem 2.17. ∎
However, just before $v_{-}$ is about to escape $K(r)$, we observe that the preimages of $D_{-}$ had boundaries nearly figure 8’s. In [3], we detected “altered” baby Julia sets which were preimage copies of a true baby Julia set with some angle identifications different from what one would see in a quadratic polynomial. Here, again there are no angle identification changes but we do have $J(r)$ consists of a tree of preimages of the quasi-disk $D_{+}$ that look one way, and then the preimages of $D_{-}$ which have the two lobes of the (near) figure 8 each mapping onto $D_{-}$, and then their preimages are “clone” conformal copies figure 8’s mapping onto figure 8’s. So there are alterations in the shape in a sense, though they’re still all quasi-disks.
Note that (in this subfamily) when $a$ is varied just enough for $v_{-}\notin K(r)$, so the assumption that $v_{-}$ maps into $D_{+}$ is a near miss—like in the right side of Figure 11—some of these statements still apply. We still have item (1), except that $D_{-}$ does not exist per se–there is still a preimage of $D_{+}$ near $v_{-}$, but it does not contain $v_{-}$. We still have item (2). Items (3) and (4) no longer apply as stated because the odd critical points escape, but consider the component that for nearby $a$ was $D_{-}$; it is still a preimage of $D_{+}$, so it looks the same, but since $v_{-}$ is not inside of it, its preimage components, rather than each containing a critical point and mapping $2:1$ onto $D_{-}$, consist of a pair of components each mapping $1:1$.
@@ 7.2. Hyperbolic components
At the beginning of Section 7 we already defined $\mathcal{H}_{2j}$ as the hyperbolic component in the bifurcation locus of $r_{n,a}$ with center $a_{2j}$ (defined in Lemma 5.1 to match the statement of Theorem 1.1) such that $r(v_{+})=v_{+}$ is a super-attracting fixed point, and $r(v_{-})=v_{+},$ and $v_{-}$ is a critical point (a non-principle root of $a^{1/2n}$).
We know the behavior at each $a_{2j}$ is different, so each $a_{2j}$ has to be in a different hyperbolic component, excluding the degenerate cases which are in the same component near $a=0$ (as was described at the beginning of Section 3): those components would correspond to $j=0,n+1$ which is why the statement of this theorem listed only $j=1,\ldots,n$.
@@
Recall from Lemma 3.1 that for $a_{2j},$ which is safely interior to $\mathcal{H}_{2j},$ $r$ is conjugate to $p_{0}(z)=z^{2}$ via a map $\phi_{a}$ which quasi-conformally maps a neighborhood of $\overline{D^{+}}$ to a neighborhood of $\overline{\mathbb{D}}$, and $\phi_{a}$ from the filled Julia set to the closed disk is conformal.
Note by Theorem 2.5, $\phi_{a}$ given by Lemma 3.1 depends continuously on $a$ for $a$ in the interior of any hyperbolic component (because the baby Julia set $D_{+}$ is always connected).
@@
remembering $v_{-}$ is a function of $a$. That is, $\Phi_{j}(a)=\phi_{a}(r_{n,a}(r_{n,a}(w)))$, where $w\in R^{-1}(v_{-})$, i.e., $w^{2n}=a$ and $w$ is any of the $n$ critical points that map to $v_{-}$. Specifically, from Lemma 2.13, at $a=a_{2j}$ the center of $\mathcal{H}_{2j}$, the $n$ critical points $w_{2l-1}=|a|^{1/2n}e^{i(\frac{\psi+2(2l-1)\pi}{2n})}$ for l=1,2,3,,n,l=1,2,3,...,n, map to $v_{-}$. Let’s just use $l=1$, remember none of these are $v_{-}$. Set $f_{n}(a)=|a|^{1/2n}e^{i(\frac{\psi+2\pi}{2n})}=|a|^{1/2n}e^{i(\frac{\psi}{2n}+\frac{\pi}{n})}=w_{1}$, so $f_{n}(a)$ is the map sending $a$ to a fixed non-principle root of $a^{1/2n},$ the one that is argument $\pi/n$ larger than $v_{+}$.
So in the hyperbolic component $\mathcal{H}_{2j},$ we have $\Phi_{j}=\phi_{a}\circ r_{n,a}\circ r_{n,a}\circ f_{n},$ where $f_{n}(a)$ is a consistent choice of $a^{1/2n},$ $r$ is a rational map in both applications a $2:1$ branched covering ramified over a critical point, and we know $\phi_{a}$ is analytic in $z$ (by Lemma 3.1, on $\overline{D_{+}(a)}$).
Now to shed light on the map $\Phi_{j}$, we’ll view it a different way. Since $a_{2j}$ is the dynamical center of a hyperbolic component (where $v_{-}$ maps to the fixed critical value $v_{+}$), let $\hat{a}_{2j}$ be a (very close) nearby point in the same component, thus its map has no non-persistent critical orbit relation (since $r_{n,\hat{a}_{2j}}(v_{-}(\hat{a}_{2j}))\neq v_{+}(\hat{a}_{2j})$). Thus, $r_{n,\hat{a}_{2j}}$ is topologically stable, so by McMullen and Sullivan [24] (as described in Section 2), we can use this as a basepoint to define a holomorphic motion of the entire sphere, in the full hyperbolic component minus its dynamical center (as there are no obstructions so this holomorphic motion extends to the punctured hyperbolic component), which respects (conjugates) the dynamics; i.e, let $\mathcal{H}_{2j}^{*}:=\mathcal{H}_{2j}\setminus\{a_{2j}\}$ and let $\tau^{2j}:\mathcal{H}_{2j}^{*}\times\hat{\mathbb{C}}\to\hat{\mathbb{C}}$ be this holomorphic motion, with notation $\tau^{2j}(a,z)=\tau^{{2j}}_{a}(z)$, and such that $\tau^{2j}({\hat{a}_{2j}},\cdot)=id,$ and
@@
Now, since $\tau$ is a holomorphic motion, for each fixed $a\in\mathcal{H}_{2j}^{*}$, $\tau^{2j}_{a}$ is an injective function of $z$ (again identity at the basepoint), and for each fixed $z$ in the dynamical plane of the basepoint map, $\tau^{2j}(\cdot,z)$ is a holomorphic function of $a$.
It turns out that due to uniqueness from conjugacy with the dynamics, we can rewrite the definition of $\Phi_{j}$ using $\tau$. We have $\tau^{2j}:\mathcal{H}_{2j}^{*}\times\hat{\mathbb{C}}\to\hat{\mathbb{C}}$, consider the restriction to $\mathcal{H}_{2j}^{*}\times\overline{D_{+}}(\hat{a}_{2j})\to\overline{D_{+}}(a)$; that is, the baby filled Julia set at the reference point $\hat{a}_{2j}$ mapping to the baby filled Julia set at the target $a$. If you fix an $a\in\mathcal{H}_{2j}^{*}$ then this restricted $\tau$ maps $\overline{D_{+}}(\hat{a}_{2j})$ to $\overline{D_{+}}(a)$. See the left half of Figure 12. Now the map $\phi_{a}$ maps $D_{+}(a)$ to $\mathbb{D}$, so if you then apply $\phi_{a}$ you again get the topological disk $\overline{\mathbb{D}}$. Instead of as above in the definition of $\Phi$ just thinking about where each different $r(v_{-})$ is in its own $D_{+}(a)$, before it is mapped into $\mathbb{D}$ by $\phi_{a}$ (which is related to where $v_{-}(a)$ is in $D_{-}(a)$), think about where $r_{\hat{a}_{2j}}(v_{-}(\hat{a}_{2j}))$ (or the preimage $v_{-}(\hat{a}_{2j})$), is moved to via $\tau$ into each $D_{+}(a)$ (or $D_{-}(a)$, respectively). This ends up being equivalent thinking, due to the uniqueness of conjugation with dynamics guaranteed by McMullen-Sullivan.
@@
Now if we apply $\phi_{a}$, we map to a point in $\mathbb{D}$ (aka, the Julia set of $z\mapsto z^{2}$). Since $a\neq a_{2j}$, $r_{n,a}(v_{-}(a))\neq v_{+}(a)$, so we know $\phi_{a}(r_{n,a}(v_{-}(a)))\neq 0$. So,

### v2

@@ 7. Topological Disks in the $r_{n,a}$ Parameter Space
Here, we consider $k=2j,$ for $j\in\{1,\ldots,n-1\}$; that is, $k$ even and $k\in\{2,\ldots,2n-2\}$.
First, for $a=a_{2j}$ as defined in Lemma 5.1, we have $r(v_{+})=v_{+}$ and $r(v_{-})=v_{+}$. Further, $v_{-}$ itself is a critical point, with
@@ 7.1. Dynamical plane
Now, recall from Lemma 3.1 that $D_{+}$ (or $D_{+}(a)$ if this is needed for clarity) is the Fatou component containing $v_{+}$ (aka $v_{+}(a)$) and on $\overline{D_{+}},$ the map $r$ is conjugate to $p_{0}(z)=z^{2}$ via $\phi_{a}$; not just for $a=a_{2j}$, but for all $a\in\mathcal{H}=\{a|\ r_{n,a}$ is hyperbolic }\}.
@@
We will assume for $a=a_{2j}$ that $v_{-}\notin D_{+}$; that is, $D_{-}\neq D_{+}$.
First we discuss the dynamics at $a=a_{2j}$. Now, $v_{-}$ is not a fixed point for $a=a_{2j},$ and so there are $n$ critical points which map to $v_{-}$ and are all distinct from it. From Lemma 2.13, the critical points $w_{2\ell-1}=|a|^{1/2n}e^{i(\frac{\psi+2(2\ell-1)\pi}{2n})}$ for $\ell=1,2,3,...,n,$ map to $v_{-}$, and the critical points $w_{2\ell}=|a|^{1/2n}e^{i(\frac{\psi+2(2\ell)\pi}{2n})}$ for $\ell=0,1,2,...,n-1,$ map to $v_{+},$ with $w_{0}=v_{+}$ in this case and our special $j$ satisfying $w_{2j}=v_{-}$. So, not counting $v_{+}$ itself, there are n1n-1 critical points mapping to $v_{+},$ one of which is $v_{-}$, and $n$ critical points mapping to $v_{-}$.
For each $\ell\in\{0,\ldots,2n-1\},$ let $D_{\ell}$ be the Fatou component containing the critical point $w_{\ell}$, so $D_{0}=D_{+}$ and for $a=a_{2j}$, $D_{2j}=D_{-}.$
Thus $r$ maps each $D_{2\ell-1}$ onto $D_{-}$ as a $2:1$ branched covering ramified over the critical point. Similarly, $r$ maps each $D_{2\ell},$ including $D_{-}$, as a $2:1$ branched cover onto $D_{+}$, ramified over the critical point in $D_{2\ell}$.
See Figure 10 for an illustration of how the Fatou components containing critical points map, in these cases of $a=a_{2j}$ so that $v_{+}$ is fixed with $v_{-}$ mapping onto $v_{+}$.
@@
Next, if we consider $a$ near $a_{2j}$, of course $D_{+}$ varies with $a$, but as shown in Lemma 3.1, $v_{+}$ is safely inside of $D_{+}$. Hence for $a$ near $a_{2j},v_{-}$ is safely inside of $D_{-}$, since $r(v_{-})=v_{+}$ at $a=a_{2j}$. As $a$ varies away from $a_{2j},v_{-}$ moves toward the boundary of $D_{-}$, its image $r(v_{-})$ moves away from $v_{+}$, heading toward the boundary of $D_{+}$ as $a$ grows. Additionally, the critical point preimages of $v_{-}$, the $w_{2\ell-1}$’s, approach the boundaries of their Fatou components $D_{2\ell-1}$, or, more accurately, the boundary of the $D_{2\ell-1}$’s tend to the $w_{2\ell-1}$’s. Since $r$ is a $2:1$ ramified cover from each component $D_{2\ell-1}$ of $r^{-1}(D_{-})$ onto $D_{-}$, as $v_{-}\to\partial D_{-}$, when $v_{-}$ hits the boundary of $D_{-}$, each $D_{2\ell-1}$ devolves into a topological lemniscate, with a figure 8 boundary with the critical point $w_{2\ell-1}$ at the crossing.
As $a$ continues to move away from $a_{2j},$ just past the lemniscate phase, there are then two Fatou components that replace $D_{2\ell-1},$ neither containing a critical point, each mapping $1:1$ onto the preimage of $D_{+}$ that is no longer $D_{-}$ because $v_{-}$ has moved out of it.
@@
We apply this result here where $\overline{D_{+}}$ is the baby filled Julia set $K_{+}$, and the quadratic polynomial we are conjugate to is $P_{0}(z)=z^{2}$ with filled Julia set $\overline{\mathbb{D}}$. Then we get a set of angle assignments on each preimage of $\partial D_{+}$, which respects the dynamics of $r_{n,a}$.
@@
Let $r=r_{n,a}$ be in the subfamily given in (1), so the Julia set $J(r)=J_{r}$ contains a baby Julia set $\overline{D_{+}}$ on which $r$ is conjugate to $P_{0}(z)=z^{2}$ on its Julia set, the closed unit disk, via a map $\phi_{a}$ which is q.c. in a neighborhood of the filled Julia sets and analytic on the interior. Assume in addition that $v_{-}$ lies in a preimage of $D_{+}$.
@@
    For components $D_{2\ell}$, $\ell=0,\dots,n-1$, each containing the critical point $w_{2\ell}$, we have
    $r\Big(\Gamma|^{\partial D_{2\ell}}\big([0,\frac{1}{2})\big)\Big)=r\Big(\Gamma|^{\partial D_{2\ell}}\big([\frac{1}{2},1)\big)\Big)=\Gamma|^{\partial D_{+}}\big([0,1)\big)$. This includes $D_{+}=D_{0}$, and $D_{-}$, for some $\ell>0$.
@@
For (1), note $r(D_{2\ell})=D_{+}$ for each $\ell=0,\dots,n-1$, where each $r(D_{2\ell})=D_{+}$ is a $2:1$ mapping.
@@
Then each of the statements about how subintervals $0,\frac{1}{2})$, $\frac{1}{2},1)$, $0,\frac{1}{4})$, $\frac{1}{4},\frac{1}{2})$, $\frac{1}{2},\frac{3}{4})$, and $\frac{3}{4},1)$ map follows directly from applying Theorem 2.17. ∎
Just before $v_{-}$ is about to escape $K(r)$, we observe that the preimages of $D_{-}$ had boundaries that are nearly figure 8’s. In [3], we detected “altered” baby Julia sets which were preimage copies of a true baby Julia set with some angle identifications different from what one would see in a quadratic polynomial. Again, in our case of interest there are no angle identification changes, but we do have $J(r)$ which consists of a tree of preimages of the quasi-disk $D_{+}$ that look one way and the preimages of $D_{-}$ which have the two lobes of the (near) figure 8 each mapping onto $D_{-}$, and then their preimages are “clone” conformal copies of figure 8’s mapping onto figure 8’s. So there are alterations in the shape in a sense, though they’re still all quasi-disks.
Note that in this subfamily, when $a$ is varied just enough for $v_{-}\notin K(r)$, so the assumption that $v_{-}$ maps into $D_{+}$ is a near miss—like in the right side of Figure 11—some of these statements still apply. We still have item (1), except that $D_{-}$ does not exist per se–there is still a preimage of $D_{+}$ near $v_{-}$, but it does not contain $v_{-}$. We still have item (2). Items (3) and (4) no longer apply as stated because the odd critical points escape, but consider the component that for nearby $a$ was $D_{-}$; it is still a preimage of $D_{+}$, so it looks the same, but since $v_{-}$ is not inside of it, its preimage components, rather than each containing a critical point and mapping $2:1$ onto $D_{-}$, consist of a pair of components each mapping $1:1$.
@@ 7.2. Hyperbolic components
At the beginning of Section 7 we defined $\mathcal{H}_{2j}$ as the hyperbolic component in the bifurcation locus of $r_{n,a}$ with center $a_{2j}$ such that $r(v_{+})=v_{+}$ is a super-attracting fixed point, $r(v_{-})=v_{+},$ and $v_{-}$ is a critical point and in particular a non-principle root of $a^{1/2n}$. This definition was given in Lemma 5.1 as to match the statement of Theorem 1.1.
We know the behavior at each $a_{2j}$ is different, so each $a_{2j}$ has to be in a different hyperbolic component, excluding the degenerate cases which are in the same component near $a=0$, as described at the beginning of Section 3: those components would correspond to $j=0,n+1$ which is why the statement of this theorem listed only $j=1,\ldots,n$.
@@
Recall from Lemma 3.1 that for $a=a_{2j},$ which is safely interior to $\mathcal{H}_{2j},$ $r$ is conjugate to $p_{0}(z)=z^{2}$ via a map $\phi_{a}$ which quasi-conformally maps a neighborhood of $\overline{D^{+}}$ to a neighborhood of $\overline{\mathbb{D}}$, and $\phi_{a}$ from the filled Julia set to the closed disk is conformal.
Note by Theorem 2.5, $\phi_{a}$ given by Lemma 3.1 depends continuously on $a$ for $a$ in the interior of any hyperbolic component since the baby Julia set $D_{+}$ is always connected.
@@
remembering $v_{-}$ is a function of $a$. That is, $\Phi_{j}(a)=\phi_{a}(r_{n,a}(r_{n,a}(w)))$, where $w\in r^{-1}(v_{-})$, i.e., $w^{2n}=a$ and $w$ is any of the $n$ critical points that map to $v_{-}$. Specifically, from Lemma 2.13, at $a=a_{2j}$ the center of $\mathcal{H}_{2j}$, the $n$ critical points $w_{2l-1}=|a|^{1/2n}e^{i(\frac{\psi+2(2l-1)\pi}{2n})}$ for l=1,2,3,,n,l=1,2,3,...,n, map to $v_{-}$. For the purpose of demonstration, we use $l=1$, remembering that no value of $l$ makes this point equal to $v_{-}$. Set $f_{n}(a)=|a|^{1/2n}e^{i(\frac{\psi+2\pi}{2n})}=|a|^{1/2n}e^{i(\frac{\psi}{2n}+\frac{\pi}{n})}=w_{1}$, so $f_{n}(a)$ is the map sending $a$ to a fixed non-principle root of $a^{1/2n},$ the one that is argument $\pi/n$ larger than $v_{+}$.
So in the hyperbolic component $\mathcal{H}_{2j},$ we have $\Phi_{j}=\phi_{a}\circ r_{n,a}\circ r_{n,a}\circ f_{n},$ where $f_{n}(a)$ is a consistent choice of $a^{1/2n},$ $r$ is a rational map in both applications a $2:1$ branched covering ramified over a critical point, and we know $\phi_{a}$ is analytic in $z$ on $\overline{D_{+}(a)}$ by Lemma 3.1.
Now to shed light on the map $\Phi_{j}$, we’ll view it a different way. Since $a_{2j}$ is the dynamical center of a hyperbolic component which is where $v_{-}$ maps to the fixed critical value $v_{+}$, let $\hat{a}_{2j}$ be a point near $a_{2j}$ in the same component, so that its map has no non-persistent critical orbit relation since now we have that $r_{n,\hat{a}_{2j}}(v_{-}(\hat{a}_{2j}))\neq v_{+}(\hat{a}_{2j})$. Thus, $r_{n,\hat{a}_{2j}}$ is topologically stable, so by McMullen and Sullivan [24] (as described in Section 2), we can use this as a basepoint to define a holomorphic motion of the entire sphere, and as there are no obstructions, this holomorphic motion extends to the punctured hyperbolic component, having removed its dynamical center. This holomorphic motion respects, or, more specifically, conjugates, the dynamics; i.e, for $\mathcal{H}_{2j}^{*}:=\mathcal{H}_{2j}\setminus\{a_{2j}\}$, let $\tau^{2j}:\mathcal{H}_{2j}^{*}\times\hat{\mathbb{C}}\to\hat{\mathbb{C}}$ be this holomorphic motion, with notation $\tau^{2j}(a,z)=\tau^{{2j}}_{a}(z)$, and such that $\tau^{2j}({\hat{a}_{2j}},\cdot)=id,$ and
@@
Now, since $\tau$ is a holomorphic motion, for each fixed $a\in\mathcal{H}_{2j}^{*}$, $\tau^{2j}_{a}$ is an injective function of $z$ which again is the identity at the basepoint, and for each fixed $z$ in the dynamical plane of the basepoint map, $\tau^{2j}(\cdot,z)$ is a holomorphic function of $a$.
Due to uniqueness from conjugacy with the dynamics, we can rewrite the definition of $\Phi_{j}$ using this $\tau$. Our original map is $\tau^{2j}:\mathcal{H}_{2j}^{*}\times\hat{\mathbb{C}}\to\hat{\mathbb{C}}$, but consider the restriction to $\mathcal{H}_{2j}^{*}\times\overline{D_{+}}(\hat{a}_{2j})\to\overline{D_{+}}(a)$; that is, let us restrict to the baby filled Julia set at the reference point $\hat{a}_{2j}$ mapping to the baby filled Julia set at the target $a$. If you fix an $a\in\mathcal{H}_{2j}^{*}$ then this restricted $\tau$ maps $\overline{D_{+}}(\hat{a}_{2j})$ to $\overline{D_{+}}(a)$. See the left half of Figure 12. Now, the map $\phi_{a}$ maps $D_{+}(a)$ to $\mathbb{D}$, so if you then apply $\phi_{a}$ you again get the topological disk $\overline{\mathbb{D}}$. Instead of as above in the definition of $\Phi$ just considering where each different $r(v_{-})$ is in its own $D_{+}(a)$ before it is mapped into $\mathbb{D}$ by $\phi_{a}$, which depends on where $v_{-}(a)$ is in $D_{-}(a)$, think about where $r_{\hat{a}_{2j}}(v_{-}(\hat{a}_{2j}))$ or its preimage $v_{-}(\hat{a}_{2j})$, is moved to via $\tau$ into each $D_{+}(a)$ or $D_{-}(a)$, respectively. This ends up being equivalent thinking, due to the uniqueness of conjugation with dynamics guaranteed by McMullen-Sullivan.
@@
Now if we apply $\phi_{a}$, we map to a point in $\mathbb{D}$, as it is the Julia set of $z\mapsto z^{2}$. Since $a\neq a_{2j}$, $r_{n,a}(v_{-}(a))\neq v_{+}(a)$, so we know $\phi_{a}(r_{n,a}(v_{-}(a)))\neq 0$. So,